### v1

# Closed Loop Interactive Embodied Reasoning for Robot Manipulation

Abstract
Embodied reasoning systems integrate robotic hardware and cognitive processes to perform complex tasks typically in response to a natural language query about a specific physical environment. This usually involves changing the belief about the scene or physically interacting and changing the scene (e.g. Sort the objects from lightest to heaviest). In order to facilitate the development of such systems we introduce a new simulating environment that makes use of MuJoCo physics engine and high-quality renderer Blender to provide realistic visual observations that are also accurate to the physical state of the scene. Together with the simulator we propose a new benchmark composed of 10 classes of multi-step reasoning scenarios that require simultaneous visual and physical measurements. Finally, we develop a new modular Closed Loop Interactive Reasoning (CLIER) approach that takes into account the measurements of non-visual object properties, changes in the scene caused by external disturbances as well as uncertain outcomes of robotic actions. We extensively evaluate our reasoning approach in simulation and in the real world manipulation tasks with a success rate above 76% and 64%, respectively.
I Introduction
The research efforts in developing systems capable of high-level perception and reasoning [1, 2, 3, 4, 5, 6] have been increasing in recent years. Many robotics systems include an agent that is required to perform a task with a specified goal given visual observations or instructions in natural language [7, 8, 1, 4, 9]. Such tasks requires a simulation environment for generating data, or on-line agent training, together with an evaluation process. Existing environments [10, 11] differ in terms of physics simulation engines and the quality of visual observations. High computational load of realistic rendering forces a compromise between real-time physics calculation and visual quality of the simulation. Realistic synthetic visual data are crucial for development of robotics systems but the collection of such data is often restricted by a cumbersome setup process and real time robot operations [12]. More powerful simulators, datasets and benchmarks are needed for a closed-loop setup where changes to the world or physical measurements (e.g., weight, elasticity, occlusions, etc.) should be accounted for as directly affecting the next step in action planning.
Figure 1: An example task presenting capabilities of the proposed environment: synthetic scene generation, instruction generation, execution of symbolic actions for manipulation followed by physics calculation and realistic rendering. Symbolic actions with corresponding targets marked in the image. Best viewed zoomed.
In this paper, we emphasise the following contributions [^1] The environment, the benchmark and the reasoning approach are released on the project webpage https://michaal94.github.io/CLIER:
$\circ$ We introduce MuBlE, a novel modular environment for simulating and executing interactive manipulation tasks characterised by high-quality visual rendering and accurate physics calculation. It employs MuJoCo for physics and Blender for rendering, which we refer to with MuBlE (MuJoCo+Blender Environment). It can generate multimodal data based on scene specification and instruction templates. An example scene and task that can be executed in this environment is shown in Fig. 1. We believe to be the first work focusing on task planning in robotics manipulation while preserving accurate physics modelling between the bodies, in contrast to discrete action spaces such as in ManipulaThor [8] and ALFRED [4].
$\circ$ We propose SHOP-VRB2, a new benchmark for evaluating embodied closed loop reasoning systems for robotic manipulation. It includes 10 types of benchmarking tasks for single- and multi-step tabletop manipulations that require combined reasoning between visual observations (recognising attributes of objects and their relations) and physical measurements (acquiring properties obtainable through interaction only, such as weight or stiffness). SHOP-VRB2 extends [13] with a challenging combination of various modalities (language, vision, and manipulation) for benchmarking Visual Reasoning, VQA (e.g. [14]) (language + vision), and Embodied question answering (EQA) (e.g. [6, 15]) (language + vision + navigation) approaches.
$\circ$ We propose a closed-loop neuro-symbolic interactive reasoning approach (CLIER), that can plan, evaluate and execute a given task (e.g., Place the heaviest mug on the box). The highlight of our approach is that it can plan for perception of invisible object properties (e.g., weight or stiffness) and account for new measurements to adjust the plan. Our reasoning agent is able to recover from a wide range of manipulation failures caused by external disturbances and uncertain outcomes of manipulation actions.
$\circ$ We evaluate our proposed Closed Loop Interactive Embodied Reasoning (CLIER) within MuBlE using the proposed SHOP-VRB2 synthetic benchmark and validate it via comparative experiments between simulated and real-world YCB [16] objects.
In the remainder of this paper, we first review the existing simulators and justify the need for a new environment. We then present the proposed MuBlE and CLIER in Sec. III. Sec. IV introduces our SHOP-VRB2 benchmark. Results and discussion are presented in Sec. V.
II Related work
TABLE I: Comparison of the MuBlE’s capabilities with the related simulators for the robotic experiments. Our MuBlE can be characterised by high quality rendering with Blender, accurate physics with MuJoCo, included visible and non-visible states, various action spaces available, targeting tabletop manipulation, modular and easily extendable.
Simulator Rendering Physics Motion planner Non-kin. states Non-visible states Action space Speed Scale Open-source
library material quality library supports
AI2-THOR [10] Unity textures ++ Unity rigid dyn./animation $\checkmark$ Discrete + room $\checkmark$
ManipulaTHOR [8] Unity textures ++ Unity AI2THOR+manip. Discrete + room $\checkmark$
ThreeDWorld [17] Unity/V-Ray (off.) Configurable $+$ Unity+FLEX rigid/particle dyn. Continuous ++ house
Robosuite [18] MuJoCo/NVISII Configurable ++ MuJoCo rigid/artic. dyn. $\checkmark$ Continuous ++ tabletop $\checkmark$
ISAACSim [11] Omniverse RTX Configurable ++ PhysX rigid/artic. dyn. $\checkmark$ (Disc.)&Cont. ++ house
Habitat 2.0 [9] Magnum 3D scans/PBR + PyBullet rigid/artic. dyn. Continuous $+$ house $\checkmark$
CoppeliaSim [19] OpenGL Gouraoud shad. + PyBullet rigid/artic. dyn. $\checkmark$ (Disc.)&Cont. ++ room
iGibson [20] PyRender/OpenGL PBR shad. + PyBullet rigid/artic. dyn. $\checkmark$ $\checkmark$ Continuous ++ house $\checkmark$
MuBle (ours) MuJoCo/Blender procedural $+$ MuJoCo rigid/artic. dyn. $\checkmark$ $\checkmark$ $\checkmark$ Disc.& Cont. + tabletop $\checkmark$
Simulation environments for manipulation focused on combined visual and language robotic tasks. State-of-the-art simulators ManipulaTHOR [8] (extension of AI2-THOR [10]), CoppeliaSim [19], or iGibson [20] provide non-photorealistic rendering, lacking accurate shades and reflections. This is in contrast to Blender that can generate photorealistic scenes with procedural object models. ISAACSim [11] is a resource-intensive, closed-source simulator with limited customisation. ThreeDWorld [17] also makes use of a commercial renderer and does not provide the infrastructure for robotics manipulation tasks, including motion planners. Unlike prior works, we pay particular attention to conditional action planning and execution in robotics manipulation while preserving accurate physics modelling between the bodies, and offer both continuous and discrete action spaces, in contrast to discrete action spaces such as in ThreeDWorld [17] and ManipulaThor [8], or pure continuous control of the robot as seen in Robosuite [18]. For example, ManipulaThor [8] considers grasping as a discrete action, while we enable to execute sequences of various custom primitive actions (e.g., approach, gripper closing, lifting, etc.) or continuous control of the robot. From these simulators, only iGibson offers non-kinematic continuous and discrete object states e.g. temperature or burned. Our MuBlE environment differs from the aforementioned manipulator simulators by a unique and open-source combination of 1) high-quality visual data and realistic physics modelling, which is a key advantage when attempting sim2real transfers (as shown in our experiments), 2) ability to select between operating on continuous action space or usage of primitive actions, and 3) considering both visible and non-visible continuous object states (e.g., weight, or stiffness). We present a comparison between different simulators in Table I.
Benchmarks for language-conditioned manipulation. The advantages of MuBlE environment enabled us to prepare a unique benchmark for closed-loop embodied reasoning (SHOP-VRB2) that extends SHOP-VRB [13]. SHOP-VRB is a dataset closely related to the robotic manipulation scenarios considered in this work. It features various objects suitable for robotic manipulation that can be easily obtained. There are many high quality 3D models of various instances of these objects which make them suitable for generating synthetic scenes via Blender rendering that includes procedurally generated materials rather than simply texturing the mesh. We therefore extend it to a dataset and a benchmark SHOP-VRB2 that enables development and testing of systems capable of reasoning on non visual attributes and performing interactive tasks.Compared to other benchmarks, such as ALFRED [4] (based on RoboTHOR), a benchmark for understanding natural language instructions in robotics, VLMbench [21], based on RLbench [22] and CoppeliaSim [19], that adds linguistic commands and realises automatic complex tasks builders, IKEA Furniture Assembly [23], that provides a benchmark for various assembly tasks and CALVIN benchmark [24] offering natural language instructions for long-horizon manipulation tasks, our SHOP-VRB2 benchmark requires long-horizon closed-loop embodied reasoning, that involves manipulation of the world to acquire knowledge about non-visual object properties in order to complete the task or answer a query.
Symbolic approaches gained interest in visual reasoning tasks and several deep learning systems incorporate neuro-symbolic programs. CLEVR-IEP [25] suggested predicting symbolic programs to be executed on disentangled scene representation, and was followed by NS-VQA [26], and NS-CL [27]. V2A [28] introduced a symbolic approach for robotic tasks using only the initial state of the scene, and was evaluated in a real robotic environment [6], however, only open-loop single-step tasks were demonstrated. Symbolic approaches were also considered in Task and Motion Planning systems (TAMP) that were often based on the formal language PDDL [29]. Classic TAMP approaches rely on predefined rules and known dynamic models [30, 31, 32] but learning based methods are increasingly replacing handcrafted heuristics in TAMP methods [33, 34, 35]. PDDLStream [36] extends PDDL to add any sub-symbolic model in a black-box way, which also opens it to the use of learning-based methods. However, it works only on a symbolic level and does not close the loop with the real world. Deep Visual Reasoning [37] directly predicts task plans by considering initial observation only. Regression Planning Networks (RPN) [38] propose a task planning model which performs regression in symbolic space using neural networks. RPN are expanded in [39] by utilising regression planning on scene graph representations to estimate the next subgoal based on the current scene and a given scene graph, i.e. solving only subpart of the overall task that our method deals with. The interaction with the real world that follows a linguistic query (e.g.measure the weight of the mug, find the sofa) is explored in embodied question answering approaches [15, 6, 40]. In [6], language-conditioned visual reasoning is combined with manipulation. Navigation of agents based on vision and language was investigated in VLN [1]. However, both approaches considered tasks that do not react to incoming measurements, i.e. reasoning is not conditioned on new observations.
Figure 2: A diagram of CLIER reasoning and MuBlE environment including interaction between the two, and use of SHOP-VRB2 benchmark. Symbols for transferred data: $\mathtt{T}$ text of the query, $\mathtt{\tilde{G}}$ prediction of scene graph elements, $\mathtt{G}$ current scene graph, $\mathtt{S}$ subgoal (symbolic program requiring physical measurements), $\mathtt{I}$ image, $\mathtt{P}$ physical observations, $\mathtt{C}$ control signal, $\mathtt{A}$ primitive action to take, $\mathtt{R}$ returned result.
Recently, several works utilised large-langauge models (LLMs) to generate action plans for robot manipulation tasks (eg.,  [41], [42]). While they show impressive versatility and adaptability to novel environments, they suffer from lack of explainability, verifiability, and repeatability. LLMs require careful prompt engineering and filtering of outputs before the robot can execute them. Furthermore, such approaches do not capture the task and trajectory details and do not enable fast-loop reactiveness. These problems have to be resolved to broadly deploy LLMs for embodied reasoning. Using data created with MuBlE environment, we train and evaluate a novel CLIER reasoning agent that can act on a natural language query in a real scene. Moreover, we present closed-loop reasoning experiments in simulation and real tabletop scenes. A highlight of our approach is that it plans for perception of invisible object properties, e.g. when stacking objects by weight. Moreover, unlike classical methods, our reasoning agent is executed on every key frame, which allows it to recover from a wide range of manipulation failures (e.g., unsuccessful grasping, or falling object during manipulation).
III MuBlE Environment and CLIER Reasoning
In this section, we present our MuBlE environment for simulating manipulation tasks and reasoning CLIER approach that makes use of this environment. The main modules and data flow are presented in Fig. 2.
MuBlE is built on $\mathtt{robosuite}$ [18] which is a simulation framework suitable for creating robotic environments inside the physics engine MuJoCo. We equip our environment with high quality rendering powered by Blender. It is designed for a generic tabletop scenario with a single robotic manipulator and a gripper. We use the same set of robots and grippers as $\mathtt{robosuite}$. We provide a set of object models in MuJoCo and their counterparts in Blender along with templates to create new items.
We first introduce the scene graph in Sec. III-A and action planning in Sec. III-B that is a part of CLIER. We then present two main components of the environment and reasoning approach: action (render) loop in Sec. III-C, and physics loop in Sec. III-D. Sec. III-E summarises our reasoning pipeline.
Figure 3: (Left) Examples of visual observations (selected frames) generated for the actions corresponding to the instruction: Stack the lightest of metal objects on the yellow object. Note that metal cans were initially picked up to measure their weight. Further, according to the measurement taken, the heavier of the cans was put down and the lighter one was picked up and stacked on the yellow plate. (Right) Example scenes and corresponding instructions in natural language generated with MuBlE (in the dataset, instructions left to right belong to tasks 7, 3, and 1 in Tab. II).
III-A Scene Graph
A real scene can be captured by a camera or a synthetic one can be rendered by Blender based on a detailed scene description. Scene graph $\mathtt{G}$ is generated for each key frame $\mathtt{I}$ starting from the first image of the scene. A keyframe is rendered after a motion corresponding to the primitive action is finished e.g. moving from object 1 to object 2, approaching grasping pose, closing the gripper, etc. We have chosen the keyframe based approach as it allows us to use Blender for rendering frames, which have sufficient quality for sim2real transfer, and sufficient speed to allow for recovering from failures with the closed loop approach.
To parse the scene into a set of objects with attributes $\mathtt{\tilde{G}}$ we apply an instance segmentation with Mask R-CNN [43], followed by classification of object attributes with ResNet [44] and pose regression with CosyPose [45]. Scene graph is a sequence of feature vectors corresponding to each object in the scene including whether the object is in the gripper or whether the gripper is closed. We use simple geometrical heuristics to create or update the scene graph $\mathtt{G}$ for keyframe $\mathtt{I}$, similarly to [39].
High-quality rendering of every frame can be used for training but it is too slow for real-time experiments. However, many tasks are achievable with low rate visual reasoning and fast feedback loops from other sources of information (e.g. force control). Example key frames rendered during execution of a task are presented in Fig. 3 (Left).
III-B Symbolic program generator and action planner
As the reasoning method we propose a multi-staged approach inspired by Neural Symbolic VQA approaches [13, 26]. Symbolic program generator is implemented as a Seq2Seq network [46] that given translates the natural language query $\mathtt{T}$ into a sequence of symbolic programs that are evaluated on the scene graph $\mathtt{G}$ for keyframe $\mathtt{I}$, producing a subgoal $\mathtt{S}$ for the task e.g. measuring weight. We employ a transformer as the action planner to predict the next primitive action $\mathtt{A}$ and its target, given the current scene graph $\mathtt{G}$ and subgoal $\mathtt{S}$.
To enable planning of various manipulation tasks we design a set of primitive actions $\mathtt{A}$ that can be easily extended in MuBlE:
  • $\mathtt{move}$ – moves the end effector towards a given target (e.g. another object or part of the table),
  • $\mathtt{approach}$ – positions the end effector in a grasping position with respect to the target object
  • $\mathtt{close\_gripper}$, $\mathtt{open\_gripper}$ – to grasp or release,
  • $\mathtt{lift}$, $\mathtt{lower}$ – lifts or lowers the end effector,
  • $\mathtt{weigh}$ – weighs the object in the gripper,
  • $\mathtt{squeeze}$ – squeezes the object in the gripper to measure its stiffness.
Non visual states, such as weight or stiffness are encoded in the scene graph after measurements, however the transformer is provided only a binary flag indicating whether the property was measured or not and the sorting reasoning is performed once all the measurements are completed.
III-C Action (render) loop
Scene graph and action planner make part of the action (renderer) loop in Fig. 2. Primitive action controller implements a control that can execute action $\mathtt{A}$ on target object in MuJoCo physics engine or real robot. Physics loop, discussed in more details in Sec. III-D), calculates physics and collects observations $\mathtt{P}$ (e.g. pose of end effector, physical measurement, force in the gripper, etc.). A control signal is generated in every step of the physics loop, until the path is completed. After the execution, a new key frame is captured in real-world setup or rendered in simulation.
Primitive action execution with obstacle avoidance require the positions and orientations $\mathtt{P}$ of the objects in the scene. Every primitive action, including gripper closure, generates a desired trajectory (position and orientation) $\mathtt{P}$ for the end effector. The action of approaching to the grasp position entails a set of pre-coded grasp sequences. We implement grasping of cylindrical objects (depending on their position and size), grasping by the edge, grasping by the handle etc. We use a simplified model as simulated grasping is often not reliable and inconsistent with real grasping. Whenever grasping pads on all gripper fingers are in contact with the same object, it is considered as grasped. Thereafter, the object’s pose is fixed with respect to the end effector. Further, whenever the gripper starts to open, the object is detached from the gripper body with an initial velocity of 0. More details can be found on GitHub page: https://michaal94.github.io/CLIER
III-D Physics loop
The physics are calculated with the use of MuJoCo engine which makes part of MuBlE and is used by CLIER as shown in Fig. 2. With a user defined timestep, the control signal $\mathtt{C}$ is applied to the manipulator, the corresponding forces are calculated and applied to all objects in the scene. Similarly to $\mathtt{robosuite}$, the control signal takes the form of the displacement of the end effector, i.e. position and orientation change, along with binary signal for opening and closing of the gripper. By default, the control signal is translated to desired joint forces by the Operational Space Controller [47], which can be customised. A control signal to the environment can be provided by the user directly or can be obtained from the primitive action controller, which calculates the error of current end effector pose and the planned trajectory.
Every step of applying the control $\mathtt{C}$ affects the physics of the scene and generates a set of observations $\mathtt{P}$, which include: position and orientation of the end effector, joint positions and velocities, status of the gripper closure. Observations of measured non-visual properties, such as stiffness, weight, or elasticity, are also included. Additionally, measurements related to all objects in the scene may be collected, such as positions, orientations, bounding boxes, indication whether the object is currently in the gripper. Any custom sensor can be added to MuBlE as in $\mathtt{robosuite}$.
III-E Reasoning pipeline
CLIER starts by parsing the visual scene to create or update its semantics and geometry scene graph. Simultaneously, a symbolic program in the CLEVR-IEP [25] format is extracted from a natural language instruction. The program is then executed on the scene graph to select the subgoals that have to be achieved on the scene, e.g. measuring the weight of the target when the graph node of the target’s weight is empty. Further, action planning regresses the task to a sequence of primitive actions for the current state of the scene. It reevaluates the next step at every pass of the action (render) loop. The reasoning is completed when either the answer is obtained from the program, or the state of the scene matches the target in task completion modules. The components of the pipeline are presented in Fig. 2 and an example task in Fig. 4.
Figure 4: An example task from the proposed dataset demonstrating the CLIER pipeline. Blue colour denotes execution of symbolic programs, the output targets from programs are shown in the green, yellow denotes executing sequence of primitive actions after having received a trigger from program (blue). Note that actions in yellow result in updates of the scene graph depicted in the top middle.
IV SHOP-VRB2 Dataset
We introduce SHOP-VRB2 dataset created in MuBlE for training and benchmarking. The dataset includes a set of scenes with instructions to perform various tasks (e.g. Stack metal objects from heaviest to lightest). Tasks are designed to enforce reasoning simultaneously on the visual observations (recognising attributes of objects and their relations) and continuous physical measurements, taking the feedback loop into account. Every example is accompanied with a ground truth sequence of actions for successful execution, along with visual observations and detailed scene graphs. Some examples can be found in Fig. 3 and 4. We include weight measurement as a representative example of estimating non visual object properties through manipulation due to the ease of repeatability for other researchers but we also demonstrate stiffness measurements. Other properties such as roughness are also possible but require another template for training and an implementation of the control. We first discuss the tools we provide in MuBlE that can be used to generate a scene and natural language instructions Sec. IV-A. We then introduce our SHOP-VRB2 benchmark in Sec. IV-B.
IV-A Data generation tools
Scene generator is a tool in MuBlE that can procedurally generate data for training models. We provide a template scene that is compatible with the renderer within the framework. The procedural algorithm takes a set of object models, which can be customised, and randomly places them on the tabletop. Desired properties of the objects such as colour, material, and size can also be randomised. We check for collisions between meshes, unlike previous approaches that check only bounding boxes. Afterwards, a maximal desired level of occlusion is ensured. Finally, we render an image of the scene, generating full ground truth, including segmentation masks for all the objects, the robot, and the tabletop, as well as a depth map. Fig. 3 (Right) presents some example scenes generated with MuBlE. Objects in the scene are described in Sec. IV.
Instruction generator is another tool in MuBlE that given the scene, generates natural language instructions or tasks that require reasoning and interaction. We propose a template-based algorithm where we improve upon CLEVR and SHOP-VRB, which are time-consuming and scale poorly when more adjectives are introduced in the templates. This is due to evaluating all possible combinations of describing words, and validating the descriptions with respect to the given scene. Instead, we implement an efficient rejection mechanism. We first generate all combinations of short descriptions for all scene objects. Short descriptions of objects are more natural than in [48, 13]. Next, we evaluate the descriptions using scene constraints e.g. the target of the instruction has to satisfy the constraint of being pickupable otherwise the description is rejected. The instruction targets are randomly selected from the validated descriptors and used to generate ground truth symbolic program. We automatically produce the ground truth actions by reasoning backwards with predefined pairwise relations e.g. lifting an object requires grasping it, grasping requires approaching to the grasping position, etc.). Fig. 3 presents example instructions generated for the given scenes using templates that are described in Sec. IV.
IV-B Benchmark data
SHOP-VRB2 scenes include $12000$ realistically rendered scenes split to train, validation, and test sets: $10000$:$1000$:$1000$. The scenes contain typical household objects which are easily accessible. Following [13], various instances of objects are included: baking trays, bowls, chopping boards, food containers, glasses, mugs, plates, soda cans, thermoses, and wine glasses. The scenes are generated with $4$ to $5$ objects per scene (their respective distribution is $47.3\%$ and $52.7\%$). We make sure the randomization of the materials (plastic, metal, glass, rubber, wood) is realistic. Out of $54317$ placed objects, a subset was selected randomly, with $1330$ instances of the least common (big chopping board model due to many possible collisions), and $3486$ instances of the most common model (bowl – less prone to collisions).
YCB scenes include 30 simulated benchmarking scenes with 9 YCB-Video [49] objects and 3 randomly generated scenes for each of the benchmarking tasks described below. 9 YCB objects were selected so that the objects share various types of visual/physical attributes (cleanser, mustard, mug, bowl, tomato can, Cheez-it, sugar box, meat can, foam brick).
TABLE II: Instruction templates corresponding to the benchmarking tasks in the proposed dataset.
No. Instruction Templates
1. Measure the weight of the $\mathtt{OBJ1}$.
2. What is the weight of all $\mathtt{OBJ1}$s?
3. Pick up the $\mathtt{WS1}$ of all $\mathtt{OBJ1}$s.
4. Place the $\mathtt{OBJ1}$ on the $\mathtt{TP1}$ part of the table.
5. Remove all $\mathtt{OBJ1}$s from the $\mathtt{TP1}$ part of the table.
6. Place the $\mathtt{WS1}$ of all $\mathtt{OBJ1}$s on the $\mathtt{TP1}$ part of the table.
7. Stack the $\mathtt{OBJ1}$ on top of the $\mathtt{OBJ2}$.
8. Place the $\mathtt{WS1}$ of all $\mathtt{OBJ1}$s on top of the $\mathtt{OBJ2}$.
9. Stack the $\mathtt{OBJ1}$ on top of the $\mathtt{OBJ2}$ on top of the $\mathtt{OBJ3}$.
10. Stack all $\mathtt{OBJ1}$s from heaviest to lightest.
Benchmarking tasks We assign one instruction to each scene to introduce more diversity in visual observations. We designed 10 classes of benchmarking tasks revolving around moving and stacking objects based on their visual (colour, material, shape), and physical properties (weight). Example templates are presented in Tab. II. The task types are closely related to the instruction templates in Tab. II and involve 1) measuring weight of a single object, 2) measuring weight of multiple objects, 3) picking up based on weight, 4) moving single object 5) moving multiple objects 6) moving based on weight 7) stacking objects 8) stacking objects according to weight 9) stacking 3 objects 10) ordering objects according to their weight. $\mathtt{OBJx}$ refers to a description of an object consisting of a set of visual properties (chosen randomly and validated), e.g. the red object presented in Fig. 1. Note that tasks may refer to either one, specific unique object, or a set of objects sharing a certain property e.g. template 1 and 2 in Tab. II. Further, $\mathtt{TPx}$ specifies a part of the table e.g. the left and the right part. Finally, $\mathtt{WSx}$ refers to weight specifier, i.e. distinction whether the lightest or the heaviest object from the set is the target of the instruction. The length of the instructions ranges between $5$ and $16$ words. The resulting sequences contain between $5$ (measure the weight of the single object) and $46$ (stacking several items according to weight) primitive actions. All instructions are accompanied with symbolic programs in CLEVR-IEP [25] format and task specifications with respect to the scene graph.
Ground truth for visual and physical observations are provided for every scene and its benchmarking instruction. An observation is taken after every primitive action is executed and contains: position and orientation of end effector, gripper status, weight measurement, ground truth scene graph, segmentation masks, and action with its target.
V Benchmarking and results
In this section we provide benchmarking results for CLIER. We show results for two datasets: 1) in simulation for SHOP-VRB2 dataset (Sec. V-A), and 2) on a set of 30 benchmarking scenes in simulation and in real-world with YCB objects (Sec. V-B).
TABLE III: (Left) Success rates for CLIER method on SHOP-VRB2 (VRB, sim) and YCB dataset (sim/real). (Right) Execution outcomes on SHOP-VRB2, incl. successes (bold) and type of failures. See the webpage linked in Sec. I for more details and video illustration.
[…]
Task type Sim Sim Real Exit code Sim
Weight single $74.0$ $66.7$ 88.9 Correct answer $13.9$
Weight multi $65.0$ $100$ 66.7 Task success $30.0$
Pick up weight $49.0$ $100$ 88.9 Task failure $0.1$
Move single $76.0$ $66.7$ 100 Execution err $14.4$
Move multi $47.0$ $100$ 44.4 Loop detected $10.8$
Move weight $23.0$ $100$ 100 Physics err $3.5$
Stack $56.0$ $66.7$ 66.7 Program err $4.5$
Stack weight $31.0$ $33.3$ 22.2 Recognition err $9.6$
Stack three $0.0$ $66.7$ 0.0 Output error $0.6$
Order weight $18.0$ $66.7$ 100 Scene inconsistent $12.6$
Overall $43.9$ $76.7$ $64.4$ Total $100\%$
Benchmarking metrics are the rate of successful task execution (e.g. stacking) and correct question answer (e.g. weight query). Additionally, we provide accuracy split into task types presented in Tab. II. Finally, for the inference tool of CLIER we provide a classification of incorrect attempts, including errors of execution, incorrect scene recognition, loop detection, timeout, inconsistency in object tracking.
The latency of various modules is the following. CosyPose inference is  0.8s, attributes recognition:  0.15s, transformer action planner:  0.02s (as run on 2080Ti). The prediction of the next action takes around 1s as measured on the hardware. These are reasonable delays given that these modules are deployed at keyframes.
V-A SHOP-VRB2 experiments
Results for CLIER are presented in Table III(Left) which shows success rates for tasks. We observe high success rates for the tasks that include manipulation of single objects (weighing or moving one object), and strong decrease in accuracy for multi-object manipulation (stacking more objects, preceding manipulation with weighting). Further, we identify the most common reasons for failure in Table III(Right) which reports the distribution of error codes from the environment. Execution error ($14.4\%$) accounts for failures in execution of primitive actions which may arise from inaccurate scene description (e.g. typically object pose). Scene inconsistency ($12.6\%$) refers to mistakes in tracking objects IDs between frames. Loop detection ($10.8\%$) arises when primitive action chains are repeated (e.g. when approaching object to grasp with misaligned position). Note that this experiment uses simple, ResNet based pose estimation in contrast to more accurate CosyPose [45] that was trained for YCB objects but not for SHOP-VRB2 objects. Finally, we believe that overall accuracy of $43.9\%$ on SHOP-VRB2 indicates that this data forms a challenging benchmark for evaluating visual and interactive reasoning.
V-B Real-world experiments
In this section, we report comparative results for real YCB scenes that mirror 30 YCB scenes introduced in Sec. IV-B.
Experimental Setup includes Franka Emika Panda [50] arm with $7$ degrees of freedom and a $2$-finger parallel gripper. An (extrinsically calibrated) Intel Realsense D455 camera is set to face the robot and capture the front view of the scene (see Fig. 5). To allow CLIER to control the real robot, we developed a thin ZMQ [51] based communication layer that emulates the same interface as for the simulated robot (i.e., motion generation using position-based servoing). Special purpose skills include measuring stiffness via squeezing the object with different forces and measuring deformation or weight via lifting and joint torque differences. Note that the weighing has a relative error of less than ±$\pm 10\%$ in the range $0-200$ g compared to the ground truth weight. The stiffness measurement has a coefficient of variation ranging from $1.6-3.4$, which is highly repeatable. However, we were not able to obtain ground truth data for stiffness.
Sim2Real Our method uses raw RGB rendered images to train the attribute recognition module (c.f. Fig. 2). This is not further tuned between YCB sim and real experiments. The same reasoning model (c.f. Sec. III-E) is evaluated in 30 synthetic and real scenes using the same CosyPose [45] on RGB images. The reasoning modules operate on the scene graph thus are agnostic to visual input.
Figure 5: The real setup with YCB objects (left), corresponding MuJoCo simulation using estimated poses (middle), and RViz visualisation of coloured pointcloud with overlaid grey models detected by CosyPose (right).
Results We mirror experiments with the real scenes in our simulation environment. To evaluate the contribution of different modules we perform testing along with ablations by replacing them with ground truth (GT) data i.e. 1) reasoning success with GT pose and GT attributes 2) pose inference with GT attributes and GT reasoning, 3) pose inference and attribute recognition with GT reasoning, 4) full inference without using GT. The results for all these combinations are in Table IV where the success is the average across 3 runs of each experiment. Note that Seq2seq network in symbolic program generator is not included in the table as it performed with $100\%$ accuracy, therefor the error for case 1) comes from transformer action planner. We observe a very similar performance between simulation and real which validates the usefulness of high quality vision and physics in MuBlE. The performance for individual tasks with YCB scenes in simulation and real world is reported in Tab. III(Left). Sim and Real differences for YCB can be attributed mostly to CosyPose object detection and its robustness to occlusion.
Typical execution errors Our closed loop reasoning approach is able to recover from various execution errors such as dropping or moving an object (see video). Typical errors during the real world execution are caused by uncertainty in pose estimation. This leads to various cases: 1) grasp succeeds because object’s real pose is withing the margin for error. 2) a different grasp is achieved e.g. over a rim instead over diameter of an open cylindrical object. 3) the object slips out when closing the gripper. 4) the gripper collides with the object during the approach and causes a robot error (safety stop reflex). The robot can recover from all except case (4) or when the object is pushed out of the workspace in (3). A varied pose of the object in the gripper (i.e., (1) and (2)) occasionally influences subsequent tasks such as stacking or weighing. Detection of occluded objects is also prone to failures. In particular, occlusion from the gripper when grasping is an issue when updating the scene graph. We use the last estimated object pose and leverage the information of the gripper holding ’something’ to overcome these pose estimation errors.
Joint limits and singularities. Robot kinematic agnostic servoing of the end-effector can lead to degenerate robot configurations. These situations include 1) joint limits, 2) the borders of the workspace, 3) self-collisions, and 4) alignment of several joint axes. To avoid running into joint limits, we start each experiment in a good standard pose and the robot always rotates to a neutral down-facing end-effector orientation before moving via the smallest rotation to the goal rotation. Self collisions are avoided by the Panda controller and are counted as robot errors.
TABLE IV: Success rate for different setups (inference vs. ground truth) in simulation and real world. C - CosyPose, A - attributes recognition, R - reasoning, ✓- module used, ✗- module substituted with ground truth information.
C A R Simulation Real
$86.7\%$ $86.7\%$
$90.0\%$ $80.0\%$
$90.0\%$ $76.7\%$
$76.7\%$ $64.4\%$
VI Discussion and Conclusions
In this paper, we presented our closed-loop approach for visual and interactive reasoning for robotic manipulation tasks: closed loop interactive embodied reasoning (CLIER). This approach is using our novel MuBlE environment that incorporates MuJoCo physics simulation with high-quality renderer and enables generation of multi-modal demonstration data for robotic manipulation tasks. This fully modular environment enables both data generation and benchmarking of simultaneous reasoning in visual and physical space. CLIER is able to incorporate observations of both visual and physical attributes of the manipulated objects into long-term reasoning. Capturing data from visual and physical measurements in the shared scene graph enables the symbolic reasoning approach and simplifies the implementation of the closed loop approach via keyframe based reasoning. The results from simulated and real-world experiments showed its ability to successfully transfer between the simulated and real environment and to recover from various errors or changes in the scene.
We believe this work provides a new challenge and a benchmark for future systems that aim at bridging the gap between physics simulation for robotic manipulators, realistic visual simulations and execution in real environments.

### v2

@@
 Abstract
-Embodied reasoning systems integrate robotic hardware and cognitive processes to perform complex tasks typically in response to a natural language query about a specific physical environment.
+Embodied reasoning systems integrate robotic hardware and cognitive processes to perform complex tasks, typically in response to a natural language query about a specific physical environment.
 This usually involves changing the belief about the scene or physically interacting and changing the scene (e.g.
 sort the objects from lightest to heaviest).
-In order to facilitate the development of such systems we introduce a new simulating environment that makes use of MuJoCo physics engine and high-quality renderer Blender to provide realistic visual observations that are also accurate to the physical state of the scene.
-Together with the simulator we propose a new benchmark composed of 10 classes of multi-step reasoning scenarios that require simultaneous visual and physical measurements.
-Finally, we develop a new modular Closed Loop Interactive Reasoning (CLIER) approach that takes into account the measurements of non-visual object properties, changes in the scene caused by external disturbances as well as uncertain outcomes of robotic actions.
-We extensively evaluate our reasoning approach in simulation and in the real world manipulation tasks with a success rate above 76% and 64%, respectively.
+In order to facilitate the development of such systems we introduce a new modular Closed Loop Interactive Embodied Reasoning (CLIER) approach that takes into account the measurements of non-visual object properties, changes in the scene caused by external disturbances as well as uncertain outcomes of robotic actions.
+CLIER performs multi-modal reasoning and action planning and generates a sequence of primitive actions that can be executed by a robot manipulator.
+Our method operates in a closed loop, responding to changes in the environment.
+Our approach is developed with the use of MuBle simulation environment and tested in 10 interactive benchmark scenarios.
+We extensively evaluate our reasoning approach in simulation and in real-world manipulation tasks with a success rate above 76% and 64%, respectively.
 I Introduction
-The research efforts in developing systems capable of high-level perception and reasoning [1, 2, 3, 4, 5, 6] have been increasing in recent years.
-Many robotics systems include an agent that is required to perform a task with a specified goal given visual observations or instructions in natural language [7, 8, 1, 4, 9].
-Such tasks requires a simulation environment for generating data, or on-line agent training, together with an evaluation process.
-Existing environments [10, 11] differ in terms of physics simulation engines and the quality of visual observations.
-High computational load of realistic rendering forces a compromise between real-time physics calculation and visual quality of the simulation.
-Realistic synthetic visual data are crucial for development of robotics systems but the collection of such data is often restricted by a cumbersome setup process and real time robot operations [12].
-More powerful simulators, datasets and benchmarks are needed for a closed-loop setup where changes to the world or physical measurements (e.g., weight, elasticity, occlusions, etc.) should be accounted for as directly affecting the next step in action planning.
-Figure 1: An example task presenting capabilities of the proposed environment: synthetic scene generation, instruction generation, execution of symbolic actions for manipulation followed by physics calculation and realistic rendering.
-Symbolic actions with corresponding targets marked in the image.
-Best viewed zoomed.
-In this paper, we emphasise the following contributions [^1] The environment, the benchmark and the reasoning approach are released on the project webpage https://michaal94.github.io/CLIER:
-$\circ$ We introduce MuBlE, a novel modular environment for simulating and executing interactive manipulation tasks characterised by high-quality visual rendering and accurate physics calculation.
-It employs MuJoCo for physics and Blender for rendering, which we refer to with MuBlE (MuJoCo+Blender Environment).
-It can generate multimodal data based on scene specification and instruction templates.
-An example scene and task that can be executed in this environment is shown in Fig.
-1.
-We believe to be the first work focusing on task planning in robotics manipulation while preserving accurate physics modelling between the bodies, in contrast to discrete action spaces such as in ManipulaThor [8] and ALFRED [4].
-$\circ$ We propose SHOP-VRB2, a new benchmark for evaluating embodied closed loop reasoning systems for robotic manipulation.
-It includes 10 types of benchmarking tasks for single- and multi-step tabletop manipulations that require combined reasoning between visual observations (recognising attributes of objects and their relations) and physical measurements (acquiring properties obtainable through interaction only, such as weight or stiffness).
-SHOP-VRB2 extends [13] with a challenging combination of various modalities (language, vision, and manipulation) for benchmarking Visual Reasoning, VQA (e.g.
-[14]) (language + vision), and Embodied question answering (EQA) (e.g.
-[6, 15]) (language + vision + navigation) approaches.
-$\circ$ We propose a closed-loop neuro-symbolic interactive reasoning approach (CLIER), that can plan, evaluate and execute a given task (e.g., Place the heaviest mug on the box).
-The highlight of our approach is that it can plan for perception of invisible object properties (e.g., weight or stiffness) and account for new measurements to adjust the plan.
-Our reasoning agent is able to recover from a wide range of manipulation failures caused by external disturbances and uncertain outcomes of manipulation actions.
-$\circ$ We evaluate our proposed Closed Loop Interactive Embodied Reasoning (CLIER) within MuBlE using the proposed SHOP-VRB2 synthetic benchmark and validate it via comparative experiments between simulated and real-world YCB [16] objects.
-In the remainder of this paper, we first review the existing simulators and justify the need for a new environment.
-We then present the proposed MuBlE and CLIER in Sec.
-III.
-Sec.
-IV introduces our SHOP-VRB2 benchmark.
-Results and discussion are presented in Sec.
-V.
+The research efforts in developing systems capable of high-level perception and reasoning have increased in recent years ([1, 2, 3, 4, 5, 6, 7, 8, 9]), including models that utilize large pretrained vision-language models for action planning such as [10, 11, 12].
+However, the ability of these models to reason iteratively through complex problems is restricted by the necessity to ground the reasoning about robot policies in sensory observations and the robot state.
+Even if some of the above-mentioned reasoning agents are able to generate and execute a plan, either their execution, their planning, or both are open loop (e.g.
+[6, 8]).
+This means that they do not account for failures during the execution (open-loop execution) and cannot react to the knowledge acquired during execution, thus cannot solve long-term planning that requires to collect and maintain a set of sequential observations that change the action plan such as ’find the softest mattress’ (open-loop planning).
+Furthermore, they typically reason only over visual observations and goals, which do not require physical measurements via interaction with the environment (see [9], where exploratory actions to find physical object properties are selected using inference in a Bayesian network).
+Such methods are unable to robustly execute tasks such as passing the lightest bottle or finding the object with the coarsest surface.
+Although some recent models (e.g., [7], [13], [5]) incorporate closed-loop reasoning, they focus on short-term tasks, e.g., manipulating an object.
+Robust and reactive long-horizon embodied reasoning remains an open problem.
+Figure 1: A diagram of CLIER reasoning implemented in MuBlE environment including interaction between the two and use of SHOP-VRB2 benchmark.
+Transferred data: $\mathtt{T}$ – text of the query, $\mathtt{\tilde{G}}$ – prediction of scene graph elements, $\mathtt{G}$ – current scene graph, $\mathtt{S}$ – subgoal (symbolic program requiring physical measurements), $\mathtt{I}$ – image, $\mathtt{P}$ – physical observations, $\mathtt{C}$ – control signal, $\mathtt{A}$ – primitive action to take, $\mathtt{R}$ – returned result.
+The main contribution of this article is the proposed Closed-Loop neuro-symbolic Interactive Reasoning framework (CLIER) [^1] The code and its documentation are released on the project webpage https://michaal94.github.io/CLIER.
+that can plan, evaluate, and execute a given task (e.g., Move the heaviest metal object to the right) (see Fig.
+1).
+The highlight of our neuro-symbolic approach is that it can plan for perception of physical object properties (e.g., weight or stiffness) and account for new measurements to adjust the plan.
+CLIER processes visual inputs to identify objects and their poses, updating a scene graph before passing the information to an action planner.
+The planner selects the next robotic action, e.g., move or weigh (see Sec.III-E).
+Each action runs in a fast feedback loop (see Primitive Action Controller in Fig.1).
+We refer to all actions as primitive actions and any resulting observations—potentially from physical interactions—as measurements.
+Our reasoning agent is able to recover from a wide range of manipulation failures.
+Figure 2: An example long-horizon manipulation task from SHOP-VRB2 implemented within the MuBlE environment [14].
+A synthetic scene is rendered in Blender every keyframe, followed by the execution of the symbolic manipulation action planned by CLIER.
+We evaluate our proposed Closed Loop Interactive Embodied Reasoning (CLIER) within the MuBlE simulation environment [14] using the SHOP-VRB2 synthetic benchmark (see an example task in Fig.
+3) and validate it via comparative experiments between simulated and real-world YCB-VRB benchmark.
+Both of these datasets require combined reasoning between visual observations (recognising attributes of objects and their relations) and physical measurements (acquiring properties obtainable through interaction only, such as weight or stiffness).
 II Related work
-TABLE I: Comparison of the MuBlE’s capabilities with the related simulators for the robotic experiments.
-Our MuBlE can be characterised by high quality rendering with Blender, accurate physics with MuJoCo, included visible and non-visible states, various action spaces available, targeting tabletop manipulation, modular and easily extendable.
-Simulator Rendering Physics Motion planner Non-kin. states Non-visible states Action space Speed Scale Open-source
-library material quality library supports
-AI2-THOR [10] Unity textures ++ Unity rigid dyn./animation $\checkmark$ Discrete + room $\checkmark$
-ManipulaTHOR [8] Unity textures ++ Unity AI2THOR+manip. Discrete + room $\checkmark$
-ThreeDWorld [17] Unity/V-Ray (off.) Configurable $+$ Unity+FLEX rigid/particle dyn. Continuous ++ house
-Robosuite [18] MuJoCo/NVISII Configurable ++ MuJoCo rigid/artic. dyn. $\checkmark$ Continuous ++ tabletop $\checkmark$
-ISAACSim [11] Omniverse RTX Configurable ++ PhysX rigid/artic. dyn. $\checkmark$ (Disc.)&Cont. ++ house
-Habitat 2.0 [9] Magnum 3D scans/PBR + PyBullet rigid/artic. dyn. Continuous $+$ house $\checkmark$
-CoppeliaSim [19] OpenGL Gouraoud shad. + PyBullet rigid/artic. dyn. $\checkmark$ (Disc.)&Cont. ++ room
-iGibson [20] PyRender/OpenGL PBR shad. + PyBullet rigid/artic. dyn. $\checkmark$ $\checkmark$ Continuous ++ house $\checkmark$
-MuBle (ours) MuJoCo/Blender procedural $+$ MuJoCo rigid/artic. dyn. $\checkmark$ $\checkmark$ $\checkmark$ Disc.& Cont. + tabletop $\checkmark$
-Simulation environments for manipulation focused on combined visual and language robotic tasks.
-State-of-the-art simulators ManipulaTHOR [8] (extension of AI2-THOR [10]), CoppeliaSim [19], or iGibson [20] provide non-photorealistic rendering, lacking accurate shades and reflections.
-This is in contrast to Blender that can generate photorealistic scenes with procedural object models.
-ISAACSim [11] is a resource-intensive, closed-source simulator with limited customisation.
-ThreeDWorld [17] also makes use of a commercial renderer and does not provide the infrastructure for robotics manipulation tasks, including motion planners.
-Unlike prior works, we pay particular attention to conditional action planning and execution in robotics manipulation while preserving accurate physics modelling between the bodies, and offer both continuous and discrete action spaces, in contrast to discrete action spaces such as in ThreeDWorld [17] and ManipulaThor [8], or pure continuous control of the robot as seen in Robosuite [18].
-For example, ManipulaThor [8] considers grasping as a discrete action, while we enable to execute sequences of various custom primitive actions (e.g., approach, gripper closing, lifting, etc.) or continuous control of the robot.
-From these simulators, only iGibson offers non-kinematic continuous and discrete object states e.g.
-temperature or burned.
-Our MuBlE environment differs from the aforementioned manipulator simulators by a unique and open-source combination of 1) high-quality visual data and realistic physics modelling, which is a key advantage when attempting sim2real transfers (as shown in our experiments), 2) ability to select between operating on continuous action space or usage of primitive actions, and 3) considering both visible and non-visible continuous object states (e.g., weight, or stiffness).
-We present a comparison between different simulators in Table I.
-Benchmarks for language-conditioned manipulation.
-The advantages of MuBlE environment enabled us to prepare a unique benchmark for closed-loop embodied reasoning (SHOP-VRB2) that extends SHOP-VRB [13].
-SHOP-VRB is a dataset closely related to the robotic manipulation scenarios considered in this work.
-It features various objects suitable for robotic manipulation that can be easily obtained.
-There are many high quality 3D models of various instances of these objects which make them suitable for generating synthetic scenes via Blender rendering that includes procedurally generated materials rather than simply texturing the mesh.
-We therefore extend it to a dataset and a benchmark SHOP-VRB2 that enables development and testing of systems capable of reasoning on non visual attributes and performing interactive tasks.Compared to other benchmarks, such as ALFRED [4] (based on RoboTHOR), a benchmark for understanding natural language instructions in robotics, VLMbench [21], based on RLbench [22] and CoppeliaSim [19], that adds linguistic commands and realises automatic complex tasks builders, IKEA Furniture Assembly [23], that provides a benchmark for various assembly tasks and CALVIN benchmark [24] offering natural language instructions for long-horizon manipulation tasks, our SHOP-VRB2 benchmark requires long-horizon closed-loop embodied reasoning, that involves manipulation of the world to acquire knowledge about non-visual object properties in order to complete the task or answer a query.
-Symbolic approaches gained interest in visual reasoning tasks and several deep learning systems incorporate neuro-symbolic programs.
-CLEVR-IEP [25] suggested predicting symbolic programs to be executed on disentangled scene representation, and was followed by NS-VQA [26], and NS-CL [27].
-V2A [28] introduced a symbolic approach for robotic tasks using only the initial state of the scene, and was evaluated in a real robotic environment [6], however, only open-loop single-step tasks were demonstrated.
+Symbolic approaches to robot planning date back to the STRIPS planner [15], but there has been renewed interest in the possibility of incorporating unprocessed sensory, in particular visual, information.
+Several deep learning systems incorporate neuro-symbolic programs.
+CLEVR-IEP [16] suggested predicting symbolic programs to be executed on a disentangled scene representation, followed by NS-VQA [17] and NS-CL [18].
+V2A [19] introduced a symbolic approach for robotic tasks using only the initial state of the scene and was evaluated in a real robotic environment [6], however, only open-loop single-step tasks were demonstrated.
 Symbolic approaches were also considered in Task and Motion Planning systems (TAMP) that were often based on the formal language PDDL [20].
-Classic TAMP approaches rely on predefined rules and known dynamic models [30, 31, 32] but learning based methods are increasingly replacing handcrafted heuristics in TAMP methods [33, 34, 35].
+Classic TAMP approaches rely on predefined rules and known dynamic models [21, 22, 23], but learning-based methods are increasingly replacing handcrafted heuristics in TAMP methods [24, 25, 26].
 PDDLStream [27] extends PDDL to add any sub-symbolic model in a black-box way, which also opens it to the use of learning-based methods.
-However, it works only on a symbolic level and does not close the loop with the real world.
+In [5], it was used for online replanning of short-horizon robotic tasks in the real world.
 Deep Visual Reasoning [28] directly predicts task plans by considering initial observation only.
-Regression Planning Networks (RPN) [38] propose a task planning model which performs regression in symbolic space using neural networks.
-RPN are expanded in [39] by utilising regression planning on scene graph representations to estimate the next subgoal based on the current scene and a given scene graph, i.e.
-solving only subpart of the overall task that our method deals with.
-The interaction with the real world that follows a linguistic query (e.g.measure the weight of the mug, find the sofa) is explored in embodied question answering approaches [15, 6, 40].
+Regression Planning Networks (RPN) [29] propose a task planning model performing regression in symbolic space using neural networks.
+RPN are expanded in [30] by utilising regression planning on scene graph representations to estimate the next subgoal based on the current scene and the scene graph, i.e.
+solving only a subpart of the overall task that our method deals with.
+The interaction with the real world that follows a linguistic query (e.g.measure the weight of the mug; find the sofa) is explored in embodied question answering approaches [31, 6, 32].
 In [6], language-conditioned visual reasoning is combined with manipulation.
-Navigation of agents based on vision and language was investigated in VLN [1].
+Navigation of agents based on vision and language was investigated in VLN [1] or [33].
 However, both approaches considered tasks that do not react to incoming measurements, i.e.
-reasoning is not conditioned on new observations.
-Figure 2: A diagram of CLIER reasoning and MuBlE environment including interaction between the two, and use of SHOP-VRB2 benchmark.
-Symbols for transferred data: $\mathtt{T}$ text of the query, $\mathtt{\tilde{G}}$ prediction of scene graph elements, $\mathtt{G}$ current scene graph, $\mathtt{S}$ subgoal (symbolic program requiring physical measurements), $\mathtt{I}$ image, $\mathtt{P}$ physical observations, $\mathtt{C}$ control signal, $\mathtt{A}$ primitive action to take, $\mathtt{R}$ returned result.
-Recently, several works utilised large-langauge models (LLMs) to generate action plans for robot manipulation tasks (eg.,  [41], [42]).
-While they show impressive versatility and adaptability to novel environments, they suffer from lack of explainability, verifiability, and repeatability.
-LLMs require careful prompt engineering and filtering of outputs before the robot can execute them.
+reasoning not conditioned on new observations.
+Recently, several works utilised large-language models (LLMs) to generate action plans for robot manipulation tasks (e.g.
+[34], [35], [36], [12], [37]).
+While they show impressive versatility and adaptability to novel environments, they suffer from a lack of fast-loop responsiveness, verifiability, and repeatability.
+LLMs require careful prompt engineering and filtering of outputs before the execution (e.g.
+[11]).
 Furthermore, such approaches do not capture the task and trajectory details and do not enable fast-loop reactiveness.
-These problems have to be resolved to broadly deploy LLMs for embodied reasoning.
+These problems have to be resolved before broader LLMs deployment.
 Using data created with MuBlE environment, we train and evaluate a novel CLIER reasoning agent that can act on a natural language query in a real scene.
 Moreover, we present closed-loop reasoning experiments in simulation and real tabletop scenes.
-A highlight of our approach is that it plans for perception of invisible object properties, e.g.
+A highlight of our approach is that it plans for the perception of invisible object properties, e.g.
 when stacking objects by weight.
-Moreover, unlike classical methods, our reasoning agent is executed on every key frame, which allows it to recover from a wide range of manipulation failures (e.g., unsuccessful grasping, or falling object during manipulation).
-III MuBlE Environment and CLIER Reasoning
-In this section, we present our MuBlE environment for simulating manipulation tasks and reasoning CLIER approach that makes use of this environment.
-The main modules and data flow are presented in Fig.
-2.
-MuBlE is built on $\mathtt{robosuite}$ [18] which is a simulation framework suitable for creating robotic environments inside the physics engine MuJoCo.
-We equip our environment with high quality rendering powered by Blender.
-It is designed for a generic tabletop scenario with a single robotic manipulator and a gripper.
-We use the same set of robots and grippers as $\mathtt{robosuite}$.
-We provide a set of object models in MuJoCo and their counterparts in Blender along with templates to create new items.
-We first introduce the scene graph in Sec.
-III-A and action planning in Sec.
-III-B that is a part of CLIER.
-We then present two main components of the environment and reasoning approach: action (render) loop in Sec.
-III-C, and physics loop in Sec.
-III-D.
-Sec.
-III-E summarises our reasoning pipeline.
-Figure 3: (Left) Examples of visual observations (selected frames) generated for the actions corresponding to the instruction: Stack the lightest of metal objects on the yellow object.
+Moreover, unlike classical methods, our reasoning agent is executed on every keyframe, which allows it to recover from a wide range of manipulation failures (e.g., unsuccessful grasping or an object falling during manipulation).
+III The Proposed CLIER Method
+In this section, we present our closed-loop interactive reasoning approach (CLIER) for long-horizon manipulation tasks that require embodied perception.
+The main modules are presented in Fig.
+1 and an example task in Fig.
+4.
+CLIER method is implemented within the MuBlE simulation environment [14], which is introduced in Sec.
+IV-A.
+In the following we introduce the individual parts of our reasoning pipeline.
+CLIER starts by parsing the visual scene (see Sec.
+III-A) to create or update its semantic and geometric scene graph (see Sec.
+III-B).
+As the reasoning agent we propose a multi-staged approach inspired by Neural Symbolic VQA approaches [38, 17].
+Simultaneously to the scene parsing, a symbolic program is extracted from a natural language instruction (see Sec.
+III-C).
+Afterwards, the Symbolic Program Executor evaluates the symbolic program on the scene graph to select the subgoals that have to be achieved on the scene, e.g.
+measuring the weight of the target when the graph node of the target’s weight is empty (see Sec.
+III-D).
+Such a subgoal is forwarded to the Action Planner, which takes into account the current state of the scene graph and the given subgoal, and selects a sequence of primitive actions to be executed in the environment.
+Each primitive action is executed within the physics loop (see Sec.
+III-E).
+After each primitive action, the Action Planner reevaluates the next step within the action (render) loop (see Sect.
+III-F).
+The reasoning is completed when the Symbolic Program Executor finishes evaluation and either the answer is obtained, or the state of the scene matches the target state in the task completion modules.
+Figure 3: (Left) Examples of simulated visual observations (selected frames) generated for the actions corresponding to the instruction: Stack the lightest of metal objects on the yellow object.
 Note that metal cans were initially picked up to measure their weight.
 Further, according to the measurement taken, the heavier of the cans was put down and the lighter one was picked up and stacked on the yellow plate.
-(Right) Example scenes and corresponding instructions in natural language generated with MuBlE (in the dataset, instructions left to right belong to tasks 7, 3, and 1 in Tab.
-II).
-III-A Scene Graph
-A real scene can be captured by a camera or a synthetic one can be rendered by Blender based on a detailed scene description.
-Scene graph $\mathtt{G}$ is generated for each key frame $\mathtt{I}$ starting from the first image of the scene.
-A keyframe is rendered after a motion corresponding to the primitive action is finished e.g.
-moving from object 1 to object 2, approaching grasping pose, closing the gripper, etc.
-We have chosen the keyframe based approach as it allows us to use Blender for rendering frames, which have sufficient quality for sim2real transfer, and sufficient speed to allow for recovering from failures with the closed loop approach.
-To parse the scene into a set of objects with attributes $\mathtt{\tilde{G}}$ we apply an instance segmentation with Mask R-CNN [43], followed by classification of object attributes with ResNet [44] and pose regression with CosyPose [45].
-Scene graph is a sequence of feature vectors corresponding to each object in the scene including whether the object is in the gripper or whether the gripper is closed.
-We use simple geometrical heuristics to create or update the scene graph $\mathtt{G}$ for keyframe $\mathtt{I}$, similarly to [39].
-High-quality rendering of every frame can be used for training but it is too slow for real-time experiments.
-However, many tasks are achievable with low rate visual reasoning and fast feedback loops from other sources of information (e.g.
-force control).
+(Right) Example simulated scenes and corresponding instructions in natural language generated with MuBlE (in the dataset, instructions left to right belong to tasks 7, 3, and 1 in Tab.
+I).
+III-A Scene Parser
+A real scene can be captured by a camera, or a synthetic scene can be rendered by Blender based on a detailed scene description within the MuBlE environment (see Sec.
+IV-A).
+We capture/render the scene in every keyframe - i.e.
+after a motion corresponding to the primitive action is finished (e.g.
+moving from object 1 to object 2, approaching grasping pose, closing the gripper, etc.).
+The keyframe-based approach allows us to use Blender for rendering frames, which have sufficient quality for sim2real transfer, and sufficient speed to allow for recovering from failures with the closed-loop approach.
 Example key frames rendered during execution of a task are presented in Fig.
 3 (Left).
-III-B Symbolic program generator and action planner
-As the reasoning method we propose a multi-staged approach inspired by Neural Symbolic VQA approaches [13, 26].
-Symbolic program generator is implemented as a Seq2Seq network [46] that given translates the natural language query $\mathtt{T}$ into a sequence of symbolic programs that are evaluated on the scene graph $\mathtt{G}$ for keyframe $\mathtt{I}$, producing a subgoal $\mathtt{S}$ for the task e.g.
-measuring weight.
-We employ a transformer as the action planner to predict the next primitive action $\mathtt{A}$ and its target, given the current scene graph $\mathtt{G}$ and subgoal $\mathtt{S}$.
-To enable planning of various manipulation tasks we design a set of primitive actions $\mathtt{A}$ that can be easily extended in MuBlE:
+To parse the scene into a set of objects with attributes $\mathtt{\tilde{G}}$, we apply segmentation with Mask R-CNN [39], followed by classification of object attributes with ResNet [40] as in SHOP-VRB [38] and pose regression with CosyPose [41].
+III-B Scene Graph
+A Scene graph $\mathtt{G}$ is generated from the parsed scene for each key frame $\mathtt{I}$ starting from the first timestep (see Fig.
+4).
+The Scene graph consists of feature vectors corresponding to each object in the scene, including categorical attributes for every object (e.g.
+material, colour), along with their positions and orientations and physical properties (e.g.
+weight, stiffness, elasticity, etc.).
+It also includes information on whether the object is in the gripper or whether the gripper is closed on it.
+Please note that the value of some of these properties might be empty (e.g.
+unknown weight of the object).
+We use geometrical heuristics to create or update the scene graph $\mathtt{G}$ for keyframe $\mathtt{I}$, similarly to [30].
+These heuristics compare the relative positions of the objects and the end effector in order to define the following attributes: whether the object is in the gripper, whether the object is within the feasible grasp, whether the object is raised above the table, and whether the gripper is placed directly above the object.
+III-C Symbolic Program Generator
+The first stage of our proposed reasoning approach inspired by neuro-symbolic VQA approaches [38, 17] is a generation of a symbolic program.
+Symbolic Program Generator is implemented as a Seq2Seq network [42] translating the natural language query $\mathtt{T}$ into a sequence of symbolic programs similarly to [38], trained with ground-truth.
+The generated program is in the CLEVR-IEP [16] format.
+Given the program sequence, it is evaluated on the current state of the scene graph $\mathtt{G}$, and either the task is finished, or a subgoal $\mathtt{S}$ necessary for task completion is produced, e.g.
+measuring the weight of the target when the graph node of the target’s weight is empty.
+When the Action Planner achieves the subgoal, the program is reevaluated and produces the next subgoal until task completion.
+III-D Action planner
+The second stage of the reasoning is an Action Planner that for every keyframe $\mathtt{I}$ and a subgoal $\mathtt{S}$, given the scene graph $\mathtt{G}$, produces a sequence of primitive actions $\mathtt{[A]}$.
+We employ a 2-layer transformer encoder as the action planner to predict the next primitive action $\mathtt{A}$ and its target, given the current scene graph $\mathtt{G}$ and the sub-goal $\mathtt{S}$, e.g.
+measuring the weight of the target when the graph node of the target’s weight is empty.
+We construct the input as the concatenation of the embeddings of the goal token and vectors describing all objects in the scene (including categorical properties, pose information, and relation to gripper based on the aforementioned heuristics).
+From this encoding, we predict the next action token, and select the target using the attention distribution over the objects training the network with ground-truth actions supervision.
+For example, when the symbolic program is queried for the task goal filter_weight[heaviest], it encounters an empty field weight in one of the scene graph nodes.
+This triggers the corresponding subgoal $\mathtt{S}$ for the Action Planner (i.e.
+measure_weight with a pointer to the graph node of the object with the empty field).
+Further, action planning regresses the task to a sequence of primitive actions for the current state of the scene.
+The Action Planner takes the current scene graph, the subgoal and its target as the input and produces the sequence of primitive actions that lead to executing the subgoal—e.g.
+$\mathtt{move}$$\mathtt{approach}$$\mathtt{close\_gripper}$$\mathtt{lift}$$\mathtt{weigh}$ for measure_weight subgoal).
+These actions are executed within the physics loop (see Sec.
+III-E).
+Non-visual states, such as weight or stiffness are encoded in the scene graph after measurements, however, the transformer is provided only a binary flag indicating whether the property was measured or not and the reasoning is performed once all the measurements are completed.
+The binary flags simplify the process as the reasoning can be done simply over the binary values instead of continuous measurement outputs.
+III-E Primitive actions and physics loop
+To enable the planning of various manipulation tasks, we designed a set of primitive actions that can be easily extended in MuBlE:
   • $\mathtt{move}$ – moves the end effector towards a given target (e.g. another object or part of the table),
@@
   • $\mathtt{weigh}$ – weighs the object in the gripper,
-  • $\mathtt{squeeze}$ – squeezes the object in the gripper to measure its stiffness.
-Non visual states, such as weight or stiffness are encoded in the scene graph after measurements, however the transformer is provided only a binary flag indicating whether the property was measured or not and the sorting reasoning is performed once all the measurements are completed.
-III-C Action (render) loop
-Scene graph and action planner make part of the action (renderer) loop in Fig.
-2.
-Primitive action controller implements a control that can execute action $\mathtt{A}$ on target object in MuJoCo physics engine or real robot.
-Physics loop, discussed in more details in Sec.
-III-D), calculates physics and collects observations $\mathtt{P}$ (e.g.
+  • $\mathtt{squeeze}$ – squeezes the object to measure its stiffness.
+A specific controller is implemented to execute each primitive action $\mathtt{A}$ on the real robot or in the MuJoCo physics engine within the MuBlE environment (see Sec.
+IV-A).
+The action of approaching to the grasp position entails a set of pre-coded grasp sequences.
+We utilize a simplified grasp simulation within MuBlE – see more details in [14].
+Each primitive action is executed within the physics loop.
+A control signal is generated at every step of the physics loop until the path is completed.
+It calculates the error for the current end effector pose and the planned trajectory and provides a control signal $\mathtt{C}$ to the environment.
+Every step of applying the control $\mathtt{C}$ affects the physics of the scene and generates a set of observations $\mathtt{P}$ (e.g.
 pose of end effector, physical measurement, force in the gripper, etc.).
+Our method assumes that for each relevant physical property, there exists at least one primitive action that produces observations of this property (e.g.
+$\mathtt{squeeze}$ for stiffness or $\mathtt{weigh}$ for mass).
+III-F Action (render) loop
+The action planner reevaluates the next step at every pass (every keyframe) of the action (render) loop that contains also scene graph generation (see Fig.
+1).
+For stability, we let the Action Planner always predict the full sequence of primitive actions until the subgoal completion.
+After the subgoal is completed (e.g., the weight field is populated in the target node of the graph), the Symbolic Program Executor resumes reasoning.
+It can either lead to producing the next subgoal (e.g., measuring the weight of another object if the query asks for comparison) or task completion.
+The Primitive Action Controller implements a control loop that can execute action $\mathtt{A}$ on target object in MuJoCo physics engine or real robot.
+The physics loop, discussed in Sec.
+III-E), calculates physics and collects observations $\mathtt{P}$ (e.g.
+pose of the end effector, physical measurement, force in the gripper, etc.).
 A control signal is generated in every step of the physics loop, until the path is completed.
-After the execution, a new key frame is captured in real-world setup or rendered in simulation.
-Primitive action execution with obstacle avoidance require the positions and orientations $\mathtt{P}$ of the objects in the scene.
-Every primitive action, including gripper closure, generates a desired trajectory (position and orientation) $\mathtt{P}$ for the end effector.
-The action of approaching to the grasp position entails a set of pre-coded grasp sequences.
-We implement grasping of cylindrical objects (depending on their position and size), grasping by the edge, grasping by the handle etc.
-We use a simplified model as simulated grasping is often not reliable and inconsistent with real grasping.
-Whenever grasping pads on all gripper fingers are in contact with the same object, it is considered as grasped.
-Thereafter, the object’s pose is fixed with respect to the end effector.
-Further, whenever the gripper starts to open, the object is detached from the gripper body with an initial velocity of 0.
-More details can be found on GitHub page: https://michaal94.github.io/CLIER
-III-D Physics loop
-The physics are calculated with the use of MuJoCo engine which makes part of MuBlE and is used by CLIER as shown in Fig.
-2.
+After the execution, a new key frame is captured in the real-world setup or rendered in simulation.
+The reasoning is completed when either the answer is obtained from the program, or the state of the scene matches the target in task completion modules (e.g.
+Weight of the object is returned or scene graph indicates that objects that were instructed to be stacked are on top of each other).
+IV Experimental setup
+IV-A MuBlE environment
+We implement and evaluate our CLIER method within the MuBlE simulations environment for manipulation tasks [14] (see Fig.
+1).
+It is built on $\mathtt{robosuite}$[43], a framework suitable for creating robotic environments in MuJoCo physics engine.
+The environment is equipped with high-quality rendering powered by Blender.
+It is designed for a tabletop scenario with a single robot manipulator and a gripper.
+It uses the same robots and grippers as $\mathtt{robosuite}$.
+It provides a set of object models in MuJoCo and their counterparts in Blender, along with templates to create new items.
+Physics calculation uses the MuJoCo engine.
 With a user defined timestep, the control signal $\mathtt{C}$ is applied to the manipulator, the corresponding forces are calculated and applied to all objects in the scene.
-Similarly to $\mathtt{robosuite}$, the control signal takes the form of the displacement of the end effector, i.e.
-position and orientation change, along with binary signal for opening and closing of the gripper.
-By default, the control signal is translated to desired joint forces by the Operational Space Controller [47], which can be customised.
-A control signal to the environment can be provided by the user directly or can be obtained from the primitive action controller, which calculates the error of current end effector pose and the planned trajectory.
-Every step of applying the control $\mathtt{C}$ affects the physics of the scene and generates a set of observations $\mathtt{P}$, which include: position and orientation of the end effector, joint positions and velocities, status of the gripper closure.
-Observations of measured non-visual properties, such as stiffness, weight, or elasticity, are also included.
-Additionally, measurements related to all objects in the scene may be collected, such as positions, orientations, bounding boxes, indication whether the object is currently in the gripper.
+The simulator provides observations including visual and physical properties of objects and end effector, including measured values for non-visual properties such as weight, stiffness, or elasticity.
 Any custom sensor can be added to MuBlE as in $\mathtt{robosuite}$.
-III-E Reasoning pipeline
-CLIER starts by parsing the visual scene to create or update its semantics and geometry scene graph.
-Simultaneously, a symbolic program in the CLEVR-IEP [25] format is extracted from a natural language instruction.
-The program is then executed on the scene graph to select the subgoals that have to be achieved on the scene, e.g.
-measuring the weight of the target when the graph node of the target’s weight is empty.
-Further, action planning regresses the task to a sequence of primitive actions for the current state of the scene.
-It reevaluates the next step at every pass of the action (render) loop.
-The reasoning is completed when either the answer is obtained from the program, or the state of the scene matches the target in task completion modules.
-The components of the pipeline are presented in Fig.
-2 and an example task in Fig.
-4.
-Figure 4: An example task from the proposed dataset demonstrating the CLIER pipeline.
-Blue colour denotes execution of symbolic programs, the output targets from programs are shown in the green, yellow denotes executing sequence of primitive actions after having received a trigger from program (blue).
+Data generation tools.
+The MuBlE environment also provides data generation tools: 1) Scene Generator places selected objects on the scene and sets their properties.
+It provides the rendered image of the scene (see Fig.
+3 (Right)) along with full ground truth data, including segmentation masks as well as a depth map.
+2) Instruction Generator given the scene, generates natural language instructions or tasks that require reasoning and interaction.
+The instruction generator also provides the symbolic program ground truth.
+Figure 4: The CLIER pipeline demonstrated on an example task from the SHOP-VRB2 dataset [14].
+Blue colour denotes a query and extracted symbolic programs, the output targets from programs are shown in the green, yellow denotes executing the sequence of primitive actions after having received a a subgoal from program (blue).
 Note that actions in yellow result in updates of the scene graph depicted in the top middle.
-IV SHOP-VRB2 Dataset
-We introduce SHOP-VRB2 dataset created in MuBlE for training and benchmarking.
-The dataset includes a set of scenes with instructions to perform various tasks (e.g.
+IV-B SHOP-VRB2 and YCB-VRB Datasets
+We evaluate the proposed CLIER method on SHOP-VRB2 and YCB-VRB benchmarking datasets [14] created in MuBlE [14] for training and benchmarking on long-horizon manipulation tasks.
+The datasets include a set of scenes with instructions to perform various tasks (e.g.
 Stack metal objects from heaviest to lightest).
-Tasks are designed to enforce reasoning simultaneously on the visual observations (recognising attributes of objects and their relations) and continuous physical measurements, taking the feedback loop into account.
-Every example is accompanied with a ground truth sequence of actions for successful execution, along with visual observations and detailed scene graphs.
-Some examples can be found in Fig.
-3 and 4.
-We include weight measurement as a representative example of estimating non visual object properties through manipulation due to the ease of repeatability for other researchers but we also demonstrate stiffness measurements.
-Other properties such as roughness are also possible but require another template for training and an implementation of the control.
-We first discuss the tools we provide in MuBlE that can be used to generate a scene and natural language instructions Sec.
-IV-A.
-We then introduce our SHOP-VRB2 benchmark in Sec.
-IV-B.
-IV-A Data generation tools
-Scene generator is a tool in MuBlE that can procedurally generate data for training models.
-We provide a template scene that is compatible with the renderer within the framework.
-The procedural algorithm takes a set of object models, which can be customised, and randomly places them on the tabletop.
-Desired properties of the objects such as colour, material, and size can also be randomised.
-We check for collisions between meshes, unlike previous approaches that check only bounding boxes.
-Afterwards, a maximal desired level of occlusion is ensured.
-Finally, we render an image of the scene, generating full ground truth, including segmentation masks for all the objects, the robot, and the tabletop, as well as a depth map.
-Fig.
-3 (Right) presents some example scenes generated with MuBlE.
-Objects in the scene are described in Sec.
-IV.
-Instruction generator is another tool in MuBlE that given the scene, generates natural language instructions or tasks that require reasoning and interaction.
-We propose a template-based algorithm where we improve upon CLEVR and SHOP-VRB, which are time-consuming and scale poorly when more adjectives are introduced in the templates.
-This is due to evaluating all possible combinations of describing words, and validating the descriptions with respect to the given scene.
-Instead, we implement an efficient rejection mechanism.
-We first generate all combinations of short descriptions for all scene objects.
-Short descriptions of objects are more natural than in [48, 13].
-Next, we evaluate the descriptions using scene constraints e.g.
-the target of the instruction has to satisfy the constraint of being pickupable otherwise the description is rejected.
-The instruction targets are randomly selected from the validated descriptors and used to generate ground truth symbolic program.
-We automatically produce the ground truth actions by reasoning backwards with predefined pairwise relations e.g.
-lifting an object requires grasping it, grasping requires approaching to the grasping position, etc.).
-Fig.
-3 presents example instructions generated for the given scenes using templates that are described in Sec.
-IV.
-IV-B Benchmark data
-SHOP-VRB2 scenes include $12000$ realistically rendered scenes split to train, validation, and test sets: $10000$:$1000$:$1000$.
-The scenes contain typical household objects which are easily accessible.
-Following [13], various instances of objects are included: baking trays, bowls, chopping boards, food containers, glasses, mugs, plates, soda cans, thermoses, and wine glasses.
-The scenes are generated with $4$ to $5$ objects per scene (their respective distribution is $47.3\%$ and $52.7\%$).
-We make sure the randomization of the materials (plastic, metal, glass, rubber, wood) is realistic.
-Out of $54317$ placed objects, a subset was selected randomly, with $1330$ instances of the least common (big chopping board model due to many possible collisions), and $3486$ instances of the most common model (bowl – less prone to collisions).
-YCB scenes include 30 simulated benchmarking scenes with 9 YCB-Video [49] objects and 3 randomly generated scenes for each of the benchmarking tasks described below.
-9 YCB objects were selected so that the objects share various types of visual/physical attributes (cleanser, mustard, mug, bowl, tomato can, Cheez-it, sugar box, meat can, foam brick).
-TABLE II: Instruction templates corresponding to the benchmarking tasks in the proposed dataset.
+Tasks are designed to enforce reasoning simultaneously on the visual observations (recognising attributes of objects and their relations) and continuous physical measurements (e.g.
+measuring invisible properties), taking the feedback loop into account.
+Every example is accompanied with a ground truth sequence of actions in CLEVR-IEP [16] format, along with visual observations and detailed scene graphs, including physical properties (see examples in Fig.
+3 and 4).
+IV-B1 Benchmarking tasks
+There are 10 classes of benchmarking tasks (see Tab.
+I; more details can be found at [14]).
+$\mathtt{OBJx}$ refers to a description of an object including visual properties, e.g.
+the red object presented in Fig.
+3 (Right).
+$\mathtt{TPx}$ specifies left/right part of the table and $\mathtt{WSx}$ is a weight specifier (lightest/heaviest).
+Instructions have $5$ to $16$ words.
+The resulting sequences contain between $5$ (measure the weight of a single object) and $46$ (stacking several items according to weight) primitive actions.
+IV-B2 SHOP-VRB2 dataset
+SHOP-VRB2 dataset [14] extends SHOP-VRB [38] dataset by introducing non-visible object properties and instructions that require closed-loop reasoning.
+SHOP-VRB2 scenes include $12000$ realistically rendered scenes with typical household objects (similarly to [38]).
+The scenes are generated as described in MuBlE [14].
+Scene generator with $4$ to $5$ objects per scene.
+SHOP-VRB2 instructions correspond to the tasks described in Sec.
+IV-B2 and are generated using MuBlE Instruction Generator tool (see Sec.
+IV-A).
+We assign one instruction to each scene.
+The instructions are randomly generated, however, they keep an equal distribution among the tasks from Table I and preserve training/test split as suggested in [14].
+TABLE I: Instruction templates and benchmarking tasks in SHOP-VRB2.
 No. Instruction Templates
@@
 10. Stack all $\mathtt{OBJ1}$s from heaviest to lightest.
-Benchmarking tasks We assign one instruction to each scene to introduce more diversity in visual observations.
-We designed 10 classes of benchmarking tasks revolving around moving and stacking objects based on their visual (colour, material, shape), and physical properties (weight).
-Example templates are presented in Tab.
-II.
-The task types are closely related to the instruction templates in Tab.
-II and involve 1) measuring weight of a single object, 2) measuring weight of multiple objects, 3) picking up based on weight, 4) moving single object 5) moving multiple objects 6) moving based on weight 7) stacking objects 8) stacking objects according to weight 9) stacking 3 objects 10) ordering objects according to their weight.
-$\mathtt{OBJx}$ refers to a description of an object consisting of a set of visual properties (chosen randomly and validated), e.g.
-the red object presented in Fig.
-1.
-Note that tasks may refer to either one, specific unique object, or a set of objects sharing a certain property e.g.
-template 1 and 2 in Tab.
-II.
-Further, $\mathtt{TPx}$ specifies a part of the table e.g.
-the left and the right part.
-Finally, $\mathtt{WSx}$ refers to weight specifier, i.e.
-distinction whether the lightest or the heaviest object from the set is the target of the instruction.
-The length of the instructions ranges between $5$ and $16$ words.
-The resulting sequences contain between $5$ (measure the weight of the single object) and $46$ (stacking several items according to weight) primitive actions.
-All instructions are accompanied with symbolic programs in CLEVR-IEP [25] format and task specifications with respect to the scene graph.
-Ground truth for visual and physical observations are provided for every scene and its benchmarking instruction.
-An observation is taken after every primitive action is executed and contains: position and orientation of end effector, gripper status, weight measurement, ground truth scene graph, segmentation masks, and action with its target.
-V Benchmarking and results
-In this section we provide benchmarking results for CLIER.
-We show results for two datasets: 1) in simulation for SHOP-VRB2 dataset (Sec.
-V-A), and 2) on a set of 30 benchmarking scenes in simulation and in real-world with YCB objects (Sec.
+IV-B3 YCB-VRB dataset
+YCB-VRB scenes include 30 simulated benchmarking scenes with 4-5 objects randomly selected from the set of 9 YCB-Video [44] objects sharing various visual/physical attributes.
+YCB-VRB Instructions are generated similarly to SHOP-VRB2.
+We generate 3 instructions per every template from Table I.
+IV-B4 Non-visual properties
+Although the SHOP-VRB2 and YCB-VRB datasets contain only weight measurement as a representative example of estimating non-visual object properties through manipulation due to the ease of repeatability for other researchers, we also demonstrate performance of our method on tasks requiring stiffness measurements (see the accompanying video [45]) to show both the ease of extending the benchmark to other physical properties as well as the applicability of the proposed CLIER method to various types of embodied actions.
+Including such a novel property only requires adding another instruction template for training and implementation of the control action to measure the given physical property.
+V Results
+In this section we provide results for CLIER: 1) in simulation for SHOP-VRB2 dataset (Sec.
+V-A), and 2) on a set of 30 benchmarking scenes in simulation and in the real world with YCB objects (Sec.
 V-B).
-TABLE III: (Left) Success rates for CLIER method on SHOP-VRB2 (VRB, sim) and YCB dataset (sim/real).
+TABLE II: (Left) Success rates for CLIER method on SHOP-VRB2 (SHOP, sim) and YCB dataset (YCB sim/real).
 (Right) Execution outcomes on SHOP-VRB2, incl.
-successes (bold) and type of failures.
+successes (bold) and types of failures.
 See the webpage linked in Sec.
@@
 Order weight $18.0$ $66.7$ 100 Scene inconsistent $12.6$
-Overall $43.9$ $76.7$ $64.4$ Total $100\%$
-Benchmarking metrics are the rate of successful task execution (e.g.
-stacking) and correct question answer (e.g.
+Overall $43.9$ $76.7$ $64.4$
+Evaluation metrics are the rate of successful task execution (e.g.
+stacking) and percentage of correctly answered questions (e.g.
 weight query).
-Additionally, we provide accuracy split into task types presented in Tab.
-II.
-Finally, for the inference tool of CLIER we provide a classification of incorrect attempts, including errors of execution, incorrect scene recognition, loop detection, timeout, inconsistency in object tracking.
+We also provide accuracy split into task types presented in Tab.
+I.
+Finally, for the inference tool of CLIER we provide a classification of incorrect attempts, including errors of execution, incorrect scene recognition, loop detection, timeout, and inconsistency in object tracking.
 The latency of various modules is the following.
@@
 V-A SHOP-VRB2 experiments
-Results for CLIER are presented in Table III(Left) which shows success rates for tasks.
-We observe high success rates for the tasks that include manipulation of single objects (weighing or moving one object), and strong decrease in accuracy for multi-object manipulation (stacking more objects, preceding manipulation with weighting).
-Further, we identify the most common reasons for failure in Table III(Right) which reports the distribution of error codes from the environment.
+Results for CLIER are presented in Table II (Left) which shows success rates for various tasks.
+We observe high success rates for the tasks that include manipulation of single objects (weighing or moving one object) and a strong decrease in accuracy for multi-object manipulation (stacking more objects, with preceding manipulation for weighting).
+Further, we identify the most common reasons for failure in Table II (Right), which reports the distribution of error codes from the environment.
 Execution error ($14.4\%$) accounts for failures in execution of primitive actions which may arise from inaccurate scene description (e.g.
@@
 Note that this experiment uses simple, ResNet based pose estimation in contrast to more accurate CosyPose [41] that was trained for YCB objects but not for SHOP-VRB2 objects.
-Finally, we believe that overall accuracy of $43.9\%$ on SHOP-VRB2 indicates that this data forms a challenging benchmark for evaluating visual and interactive reasoning.
-V-B Real-world experiments
-In this section, we report comparative results for real YCB scenes that mirror 30 YCB scenes introduced in Sec. IV-B.
-Experimental Setup includes Franka Emika Panda [50] arm with $7$ degrees of freedom and a $2$-finger parallel gripper.
+Finally, we believe that the overall accuracy of $43.9\%$ on SHOP-VRB2 indicates that this data forms a challenging benchmark for evaluating visual and interactive reasoning.
+V-B YCB-VRB and real-world experiments
+In this section, we report comparative results for real YCB scenes that mirror 30 YCB scenes introduced in Sec. IV-B2.
+Experimental Setup includes the Franka Emika Panda [46] arm with $7$ degrees of freedom and a $2$-finger parallel gripper.
 An (extrinsically calibrated) Intel Realsense D455 camera is set to face the robot and capture the front view of the scene (see Fig.
@@
 However, we were not able to obtain ground truth data for stiffness.
-Sim2Real Our method uses raw RGB rendered images to train the attribute recognition module (c.f.
-Fig.
-2).
-This is not further tuned between YCB sim and real experiments.
-The same reasoning model (c.f.
-Sec.
-III-E) is evaluated in 30 synthetic and real scenes using the same CosyPose [45] on RGB images.
-The reasoning modules operate on the scene graph thus are agnostic to visual input.
-Figure 5: The real setup with YCB objects (left), corresponding MuJoCo simulation using estimated poses (middle), and RViz visualisation of coloured pointcloud with overlaid grey models detected by CosyPose (right).
-Results We mirror experiments with the real scenes in our simulation environment.
-To evaluate the contribution of different modules we perform testing along with ablations by replacing them with ground truth (GT) data i.e.
+Sim2Real.
+Our method uses raw RGB rendered images to train the attribute recognition module (see Fig.
+1).
+This is not further tuned between the YCB sim and real experiments.
+The same reasoning model (see Sec.
+III) is evaluated in 30 synthetic and real scenes using the same CosyPose [41] on RGB images.
+The reasoning modules operate on the scene graph and thus are agnostic to visual input.
+Figure 5: The real setup with YCB objects (left), corresponding MuJoCo simulation using estimated poses (middle), and RViz visualisation of colored pointcloud with overlaid gray models detected by CosyPose (right).
+Results.
+We mirror experiments with real scenes in our simulation environment.
+To evaluate the contribution of different modules we perform testing along with ablations by replacing them with ground truth (GT) data, i.e.
 1) reasoning success with GT pose and GT attributes 2) pose inference with GT attributes and GT reasoning, 3) pose inference and attribute recognition with GT reasoning, 4) full inference without using GT.
-The results for all these combinations are in Table IV where the success is the average across 3 runs of each experiment.
-Note that Seq2seq network in symbolic program generator is not included in the table as it performed with $100\%$ accuracy, therefor the error for case 1) comes from transformer action planner.
-We observe a very similar performance between simulation and real which validates the usefulness of high quality vision and physics in MuBlE.
-The performance for individual tasks with YCB scenes in simulation and real world is reported in Tab.
-III(Left).
+The results for all these combinations are in Table III, where the success is the average across 3 runs of each experiment.
+Note that the Seq2seq network in the symbolic program generator is not included in the table as it performed with $100\%$ accuracy, therefore the error for case 1) comes from the transformer action planner.
+We observe a very similar performance between simulation and real world which validates the usefulness of high-quality vision and physics in MuBlE.
+The performance for individual tasks with YCB scenes in simulation and the real world is reported in Tab.
+II(Left).
 Sim and Real differences for YCB can be attributed mostly to CosyPose object detection and its robustness to occlusion.
-Typical execution errors Our closed loop reasoning approach is able to recover from various execution errors such as dropping or moving an object (see video).
+Typical execution errors.
+Our closed loop reasoning approach is able to recover from various execution errors such as dropping or moving an object (see video [45]).
 Typical errors during the real world execution are caused by uncertainty in pose estimation.
@@
 A varied pose of the object in the gripper (i.e., (1) and (2)) occasionally influences subsequent tasks such as stacking or weighing.
-Detection of occluded objects is also prone to failures.
+The detection of occluded objects is also prone to failures.
 In particular, occlusion from the gripper when grasping is an issue when updating the scene graph.
-We use the last estimated object pose and leverage the information of the gripper holding ’something’ to overcome these pose estimation errors.
+We use the last estimated object pose and leverage the information of the gripper holding ‘something’ to overcome these pose estimation errors.
 Joint limits and singularities.
-Robot kinematic agnostic servoing of the end-effector can lead to degenerate robot configurations.
-These situations include 1) joint limits, 2) the borders of the workspace, 3) self-collisions, and 4) alignment of several joint axes.
-To avoid running into joint limits, we start each experiment in a good standard pose and the robot always rotates to a neutral down-facing end-effector orientation before moving via the smallest rotation to the goal rotation.
+Cartesian control of the robot end-effector without taking the joint configuration into account can lead to degenerate configurations.
+These include 1) joint limits, 2) borders of the workspace, 3) self-collisions, and 4) alignment of several joint axes.
+To avoid running into joint limits, we start each experiment in a good standard pose, and the robot always rotates to a neutral down-facing end-effector orientation before moving via the smallest rotation to the goal rotation.
 Self collisions are avoided by the Panda controller and are counted as robot errors.
-TABLE IV: Success rate for different setups (inference vs.
-ground truth) in simulation and real world.
+TABLE III: Success rate for different setups (inference) in simulation and real world.
 C - CosyPose, A - attributes recognition, R - reasoning, ✓- module used, ✗- module substituted with ground truth information.
@@
 $76.7\%$ $64.4\%$
-VI Discussion and Conclusions
-In this paper, we presented our closed-loop approach for visual and interactive reasoning for robotic manipulation tasks: closed loop interactive embodied reasoning (CLIER).
-This approach is using our novel MuBlE environment that incorporates MuJoCo physics simulation with high-quality renderer and enables generation of multi-modal demonstration data for robotic manipulation tasks.
-This fully modular environment enables both data generation and benchmarking of simultaneous reasoning in visual and physical space.
+VI Conclusions
+In this paper, we presented our Closed-Loop Interactive Embodied Reasoning (CLIER) for robotic manipulation tasks.
 CLIER is able to incorporate observations of both visual and physical attributes of the manipulated objects into long-term reasoning.
-Capturing data from visual and physical measurements in the shared scene graph enables the symbolic reasoning approach and simplifies the implementation of the closed loop approach via keyframe based reasoning.
+Capturing data from visual and physical measurements in the shared scene graph enables the symbolic reasoning approach and simplifies the implementation of the closed loop approach via keyframe-based reasoning.
+The CLIER utilises MuBlE environment [14] that incorporates MuJoCo physics simulation with high-quality renderer and enables generation of multi-modal demonstration data for robotic manipulation tasks.
 The results from simulated and real-world experiments showed its ability to successfully transfer between the simulated and real environment and to recover from various errors or changes in the scene.
-We believe this work provides a new challenge and a benchmark for future systems that aim at bridging the gap between physics simulation for robotic manipulators, realistic visual simulations and execution in real environments.